OpenStudy (anonymous):

describe how to simplify each expression: 16^1/2

OpenStudy (anonymous):

\[x^{\frac{a}{b}}=\sqrt[b]{x^a}\] \[16^{\frac{1}{2}}=\sqrt{16^1}=\sqrt{16}=4\]

OpenStudy (anonymous):

\[16^{1/2}\] \[=\sqrt{16}\] \[=\sqrt{4^{2}}\] =4

OpenStudy (anonymous):

well first say raised to the half is the same thing as square root of a number. and then you realize: 16 square root is 4.